OpenStudy (anonymous):

find all solutions for 10sin(x) ^2 - sin (x) = 2

myininaya (myininaya):

\[10\sin^2(x)-\sin(x)-2=0\] let u=sinx so u^2=(sinx)^2 \[10u^2-u-2=0\] \[10u^2-5u+4u-2=0\] \[5u(2u-1)+2(2u-1)=0\] (5u+2)(2u-1)=0 u=-2/5 sinx=-2/5 u=1/2 sinx=1/2 solve these for x

OpenStudy (anonymous):

wowee! that was quick!

OpenStudy (anonymous):

you are a factoring somebody! love how you split the middle term! i can't do it

OpenStudy (anonymous):

of course you chickened out on how to solve \[\sin(x)=-\frac{2}{5}\] for x!

myininaya (myininaya):

satellite it is easy

myininaya (myininaya):

ax^2+bx+c just find two factors of a*c that add up to be b say you found m*n=a*c and m+n=b just replace bx with mx+nx and woolah!

OpenStudy (anonymous):

sinx = -2/5, so x=inverse sin(-2/5), which is in the 3rd or fourth quadrant so x = 180+23.6 degrees, or 360-23,6 degrees